OpenStudy (anonymous):

Compare and contrast, in complete sentences, the difference between the products of a2a4 and (a2)4. Provide the expanded form and product of each.

OpenStudy (anonymous):

a^2+a^2+a^2+a^2 is that the expanded form

OpenStudy (anonymous):

i just need help comparing and contrasting

OpenStudy (anonymous):

\(a^2a^4\) gives \(a^{2+4} \implies a^6\) \((a^2)^4\) gives \(a^{2 \times 4} \implies = a^8\)

OpenStudy (anonymous):

but how can i compare and contrast them in complete sentences can you help me

OpenStudy (anonymous):

i am not sure what they mean by complete sentence but waterineyes is right

OpenStudy (anonymous):

In case of multiplication with the same bases. their exponents add up.. \[\large a^m.a^n = a^{m+n}\]

OpenStudy (anonymous):

so i can say that in this case of multiplication with the same bases their exponents add up ?

OpenStudy (anonymous):

But if the bases have power of the power then the exponents get multiply: \[\large (a^m)^n = a^{m \times n}\]

OpenStudy (anonymous):

Yes you are right rebecca..

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Welcome dear...

OpenStudy (anonymous):

:)